OpenStudy (anonymous):

how would you find the slope intercept form for through (4,-1), m=0

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

I am totally lost when it comes to slope intercept. I am hoping someone will explain it to me in a way I can understand

OpenStudy (skullpatrol):



OpenStudy (anonymous):

Well, you have an m that is equal to 0 right? So: y = mx+b and you have m = 0 so: y = 0x+b so y = b. That means if y = -1 (given) b = -1 So your equation becomes: y = -1.

OpenStudy (anonymous):

oh thank you so much Zelda, that's what I needed to see to understand how to work the other problems I have like that

OpenStudy (anonymous):

Hmm, but remember this is only when m = 0. If you have say, m = 2? Then what happens is (using same point) (4,-1) Then this happens: y = mx+b --> y = 2x+b. Now that you know y and x, plug in and solve for b. Once you get b, put that into the original point slope equation.

OpenStudy (anonymous):

oh okay. so would I use this same formula if I was working on a problem for example like through (2,-4) and (0,-7)

OpenStudy (anonymous):

Yup, fist use the slope forumla, then do it the second way.

OpenStudy (anonymous):

okay im going to attempt it and maybe you can let me know if I am on the right track

OpenStudy (anonymous):

Sure thing kiddo.